OpenStudy (almightyns):

Find the product and simplify your answer -n^4(8n^3+4)

OpenStudy (studyloverhojea):

n^4(8n^3 +4) --> n^4(8n^3) + n^4(4) 8n^4 +^3 = 8n^7 8n^7 + n^4(4) = 8n^7 + 4n^4 = 12n^11 The answer is 12n^11. I think.

OpenStudy (mathstudent55):

\(-n^4(8n^3+4)\) Distribute \(-n^4\) \(= -8n^7 - 4n^4\) Since n^7 and n^4 are different variable parts, the two terms above are not like terms, and cannot be combined, so the second line above is the final answer.